OpenStudy (anonymous):

Two balloons having the same charge of 1.2 × 10-6 coulombs each are kept 5.0 × 10-1 meters apart. What is the magnitude of the repulsive force between them? (k = 9.0 × 109 newton·meter2/coulombs2) 1.2 × 10-2 newtons 3.2 × 10-3 newtons 6.5 × 10-3 newtons 3.0 × 10-4 newtons 5.2 × 10-2 newtons

OpenStudy (aaronq):

\(\sf \huge F=k\dfrac{q_1*q_2}{r^2}\)

OpenStudy (anonymous):

I dont know how to plug in everything @aaronq

OpenStudy (aaronq):

q is the charge, r is the radius, k is the constant \(\sf F= (9.0 *10^9~Nm^2/C^2 )\dfrac{(1.2*10^{-6}C)*(1.2*10^{-6}C)}{(5.0 *10^{-1}m)^2}\)

OpenStudy (anonymous):

Thank you! This really helps my mind gets jumbled with these type of questions

OpenStudy (aaronq):

no problem! they can be a bit daunting at first but you'll get the hang of it. Post physics q's in the correct section next time though :P

OpenStudy (anonymous):

This is my first time using this so Im still figuring stuff out but I will and thank you again

OpenStudy (aaronq):

okay cool. if you want to see an introductory guide or something go to the Openstudy feedback section and post "how do i use this site" or something. People are always there to help out with that. good luck with your studies!